408juju:

Given the function h(x) = 3(5)x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points) i just need answer please. ill give modded cars on gta ps4 if help

supie:

HAHA MODDED CARS ON GTA

408juju:

:(

supie:

Is this a test question?

408juju:

no its my last question on my quiz

supie:

Quiz, ok.

408juju:

need a screen shot ?

darkknight:

yes plz, send a screenshot instead

408juju:

darkknight:

average rate of change is essentially the the slope formula. (y2 - y1)/(x2 - x1)

darkknight:

for section A, from x = 0 to x = 1 plug in 0 and 1 into the equation and solve for y. Then you will have your 2 coordinates

darkknight:

From there, plug into the formula and see what you get.

408juju:

bro i got a f in math rn my dad is gonna beat me and my uncle is gonna jump in like its wwe smackdown but he likes to do some werid stuff

408juju:

i just need the answer before my dad starts hitting me and my mom being his number one fan in the background

darkknight:

... can you do what i asked you to do?

supie:

lol ikr

408juju:

okok how do i do that ?

darkknight:

Nope, you are not getting a direct answer, work with me lad

408juju:

evaluate

408juju:

?

darkknight:

so we have the equation 3(5)^x If 0 goes for x then 5^0 = 1 3 times 1 is 3 coordinate 1 = (0,3) If we plug in 1 for x 3(5)^1 = 3(5) =15 coordinate 2 = (1,15)

408juju:

bro that one hurt the heck out my head but im listening...

darkknight:

now we plug it into the slope formula y2 =15 y1 = 3 x2 = 1 x1 = 0 so now (15-3)/(1-0)

darkknight:

and 15-3 = 12 1 - 0 = 1 12 is the average rate of change

darkknight:

Can you try to find the average rate of change for section B?

408juju:

yea give me a second cause im trying to read this, understand it then look at the question for part b

darkknight:

yes ofc, if you need further clarification then ask away

408juju:

ty

408juju:

what number do i start with because you started with 3(5)^x but for part b says only x= 2 to x=3

darkknight:

you can decide, lets just assign x1 = 2 and x2 = 3 for the sake of understanding This means that the y value obtained when we plug in x = 2 into the equation will be assigned to y1. Because if x1 is plugged in, y1 is the output If x2 is plugged in y2 is the output

darkknight:

so we want to find from x = 2 to x = 3 Now lets say x1 = 2 and x2 = 3 Find y1 and y2 can you?

408juju:

bro im kinda lost sorry. so do i use x1=2 and x2=3 to find it? if so am i supposed to put it in an equation ?

darkknight:

yes its all good if your lost. So x1 =2 and x2 = 3 So we have the equation 3(5)^x plug in the value for x1 into x in the equation right above this, and the y-value you get is y1 similarly for x2, plug into the equation 3(5)^x for x and the y value you get is y2. Can you tell me what y1 and y2 are now?

408juju:

wait is y1 2 and y2 3 ??

supie:

@darkknight

darkknight:

no x1 is 2 and x2 is 3, you have to plug in 2 and 3 into the equations above to find what y1 and y2 are.

darkknight:

Okay, so x1 = 2 ill plug in 2 for x in this function 3(5)^x Now i have 3(5)^2 which is 75

408juju:

ohh

darkknight:

so do the same for x2, and find what y2 is

408juju:

so your telling me to do 3(5)^3

408juju:

am i mutiplying orr ?

darkknight:

yes, that is what you are going to do.

darkknight:

you have to evaluate 3(5)^3

408juju:

375 ?

darkknight:

yes, x2 = 3 and y2 = 375 x1 = 2 and y1 = 75. plug into this formula (y2-y1)/(x2-x1)

408juju:

ok

408juju:

like on a calculator ?

darkknight:

sure, you just have to plug in the values and find what the final value is

408juju:

ok ok

408juju:

still mutiplying ?

darkknight:

idk what you are saying, you have to evaluate

408juju:

oh my bad

darkknight:

all good, type in what you get for the average rate of change

408juju:

im putting (y2-y1)/(x2-x1) on the equation

408juju:

?*

darkknight:

Have you done it yet? you don't need to ask for confirmation every second. Just plug in values and evaluate

408juju:

lmao im trying to make sure cause i dont wanna do it wrong sorry

darkknight:

did you get your answer?

darkknight:

Did you get your answer?

supie:

@408juju The man is talkin to ya.....